OpenStudy (anonymous):

After the softball game, Alise's coach bought each of the 13 players a double-scoop or triple-scoop ice cream cone at the local ice cream shop. The coach bought a total of 33 scoops of ice cream. Use a model to determine how many girls got a triple-scoop ice cream cone.

OpenStudy (anonymous):

x+y=13 2x+3y=33

OpenStudy (anonymous):

where x is double scoop and y is triple

OpenStudy (anonymous):

7

OpenStudy (anonymous):

6 and 7

OpenStudy (anonymous):

6 is double and 7 is triple

OpenStudy (anonymous):

Solve the following for x and y. y is equal to the number of triple scoopers. {2 x + 3 y = 33, x + y = 13} x=6, y=7

OpenStudy (anonymous):

i can only choose one answer